OpenStudy (anonymous):

find the balance in the account. $500 principal earning 4% compounded quarterly, after 6 years.

OpenStudy (lexipoo1998):

I think its $620

OpenStudy (lexipoo1998):

cuz its 500 * .04=20 *6 = 120 so then i did 120+500 and got 620. Unless i misread the question or made a mistake i think that is right

OpenStudy (anonymous):

i understand how you did it but for some reason that answer isnt shown on my work. is there possibly another way of figuring it out?

OpenStudy (lexipoo1998):

Oh um..... Well what are you answer choices?

OpenStudy (anonymous):

they are A. 582.36 B. 613.58 C. 634.87

Parth (parthkohli):

\[500\left(1 + \dfrac{4}{400}\right)^{6\times 4}\]calculator.

OpenStudy (lexipoo1998):

hmmmmmmmmmm...... well I dont know how else to solve it to get one of those answers so hopefully someone else will be able to help you. Sorry.. ;/

Parth (parthkohli):

\[500 \times (1.01)^{24}\]Enter that in yer calculator.

OpenStudy (anonymous):

no worries thanks for your help. and Ah okay thank you very much parthkohli